# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

### v2

@@ Abstract
Large Language Models (LLMs) have introduced significant advancements to the capabilities of Natural Language Processing (NLP) in recent years. However, as these models continue to scale in size, memory constraints pose substantial challenge. Key and Value cache (KV cache) quantization has been well-documented as a promising solution to this limitation. In this work, we provide two novel theorems aimed at enhancing KV quantization methods. Our first theorem, termed Key-Value Norm Disparity, states that the key weight matrices by nature carry richer information compared to the value weight matrices, as evidenced by higher spectral and Frobenius norms across most of the layers. Our second theorem, Key-Driven Quantization, posits that prioritizing the quantization precision of keys over values induces significant improvements to the overall quantization performance. In particular, assigning greater precision to the keys compared to the values achieves a higher degree of precision reduction with minimal impact on model accuracy. We validate these theorems through theory and extensive experiments on several state-of-the-art LLM architectures and benchmarks. These findings offer valuable guidelines for improving KV cache quantization strategies, facilitating more efficient memory utilization without compromising model performance across diverse NLP tasks. Source code is available at https://github.com/mohsenhariri/spectral-kv.
@@ 1 Introduction
LLMs have introduced significant advancements to the capabilities of NLP via their ability to express and capture intricate ideas [1, 2], and a key reason for these advancements is model scaling. GPT has seen a tremendous increase across its four versions: GPT-1 at 117M [3], GPT-2 at 1.5B [4], GPT-3 at 175B [5], and GPT-4 at 1.8T parameters [6]. In the open-source space, Llama has grown to 2T [7], Mistral Large is now 123B [8], and DeepSeek V3 is at 671B [9].
However, with this growth in model size comes a prohibitive memory bottleneck in the KV cache [10]. An increase in parameter count requires an increase in context-window length to accommodate chain-of-thought reasoning, which in turn expands the KV cache central to the attention unit. That expansion can impose a restrictive memory bottleneck on the LLM. Already, supported context windows in SOTA systems have grown significantly [11, 12, 13, 14], underscoring the need for a solution. One powerful remedy is quantization, wherein high-precision numbers are reduced to low-precision formats [15, 16]. In LLMs, this can be applied to the KV-cache tensors (e.g., from BF16 to INT4) yielding substantial memory savings with minimal performance loss.
@@
14B 0.76 0.87
Figure 1: Keys carry greater importance. (Left): Keys exhibit higher norms than values across all Llama 3.3-70B layers. (Right): GSM8K accuracy is higher under $\mathrm{K}_{2}\mathrm{V}_{4}$ (2-bit keys, 4-bit values) than $\mathrm{K}_{4}\mathrm{V}_{2}$.
This process is known as KV cache quantization. Extensive work has been done in designing novel KV quantization frameworks, yet none so far have provided deep investigation into mixed-precision schemes, in particular those which focus on the bit allocations split between the Key and Value caches. Instead, previous works have treated the bit allocation as a hyperparameter, seeing it as an object for optimization via grid search rather than taking a deeper dive into the properties of Keys and Values. We provide further motivation for the study in the form of a full description of the current literature, which is detailed in Section 2.
Therefore, there is a present need for a deep exploration into the inherent differences between the Key and the Value caches, as well as a change to the way that quantization is performed. This study aims to fill this need by uncovering fundamental theoretical ideas to improve upon KV cache quantization methods. A preview of these results is illustrated in Figure 1 (left), showing that the spectral norms of the Keys consistently exceed those of the values. Additionally, Figure 1 (right) demonstrates that accuracy across four different models is notably higher for the K4V2 quantization configuration compared to the K2V4 configuration. Our main contributions are summarized as follows:
  • We prove that the expected values of the spectral norm and Frobenius norm are consistently higher for Key weight matrices compared to Value weight matrices (Key-Value Norm Disparity theorem), implying that Key matrices inherently carry more information. We provide detailed empirical validation through observational studies on multiple LLMs.
  • Building on the Key-Value Norm Disparity, we introduce the Key-Driven Quantization theorem, positing that assigning higher bit precision to Keys over Values significantly enhances quantization efficiency. Specifically, this approach allows for greater precision reduction while maintaining minimal accuracy loss.
  • We substantiate the Key-Driven Quantization theorem through two primary forms of evidence: (1) a theoretical analysis explicitly linking it to the previously established Key-Value Norm Disparity, and (2) comprehensive empirical evaluations involving multiple SOTA LLMs, three datasets, and two quantization frameworks.
Section 2 addresses gaps in the existing literature on KV cache quantization, followed by Section 3.1, which introduces the Key-Value Norm Disparity theorem and its supporting evidence. Section 3.2 presents the Key-Driven Quantization theorem along with its corresponding proof, and Section 4 reports the empirical evaluation of both theorems. Finally, Section 5 summarizes the main contributions and discusses their limitations.
2 Related Work
KV Quantization Background.
Quantization frameworks can be categorized in several ways. One way is the time at which the quantization is performed, which is grouped into training-time and post-training (PTQ) quantization techniques [17]. Training-time quantization frameworks can struggle with the usual costs of neural network training, including the dependence on labelled data and longer training times, but these frameworks can achieve more accurate performance due to targeting the weights or activations during the training process. In contrast, post-training quantization techniques are the reverse, struggling with sometimes less accurate performance due to dealing with the values after the training has been completed but achieving lower computation overhead and no longer relying on labelled data which may not be available [16].
There are several settings for PTQ: weight-only, weight-activation, and KV quantization. Weight-only quantization focuses on the weights and achieves strong performance but can struggle with memory overhead [18, 19, 20]. Weight-activation quantization focuses on both the weights and the activations and avoids memory overhead but can struggle with achieving high accuracy [21, 22, 23]. In contrast, KV quantization offers the memory savings seen in weight-activation methods while also maintaining the performance of weight-only quantization [24]. This is due to the KV cache memory overhead inflation caused by longer sequences [25].
KV quantization frameworks can be further categorized into three groups based on the exact quantization scheme being used for the keys and values: outlier redistribution, fixed-precision, and mixed-precision [26]. In outlier redistribution, the outliers in the cache are moved elsewhere or smoothed out. Several outlier redistribution frameworks have been built to varying levels of success [22, 20, 23]. In fixed-precision, the same number of bits is used to quantize certain parts of the cache while in mixed-precision, different numbers of bits are used. Fixed-precision frameworks have also been built to varying levels of success, though they all suffer from a similar problem of neglecting the inherent properties within the KV cache as well as the varying importance of tokens [27, 28, 29].
More focus has been devoted to the third type of KV quantization framework, mixed-precision, but as yet, these works have not looked much into asymmetric allocation into keys and values [30, 24, 31, 32, 33, 34, 35]. KVQuant, for example, focuses on defining an outlier for each vector, such as channel or token, to achieve what the authors term dense-and-sparse quantization. Only KVTuner, QAQ, and SKVQ discuss the asymmetric allocation in some detail, which is described below.
KV Quantization Gaps.
As described above, extensive work has been done in designing new KV quantization frameworks, yet none so far have conducted deeper investigation into mixed-precision schemes. In particular, no studies have focused on making a bit allocation split between the key matrix and the value matrix, nor have they provided a deep theoretical explanation for the inherent differences between the two matrices. Only three studies to date have touched upon this area, with none providing explanations for potential differences between the key and value matrix importance.
The closest study has been a KV quantization framework called KVTuner, which employs layerwise quantization based on the sensitivity of the layers to the KV cache quantization mode and precision pairs [35]. The study makes a brief mention of the difference in attention score errors and perplexity score errors, observing that precision pairs with more bits allocated for keys than for values achieves similar accuracy than for even-split pairs, such as K8V4 versus K8V8. However, the study makes no further analysis into this observation, providing no reason for why this difference is observed. In addition, the quantization approach does not use any underlying insight on this difference between key and values, instead opting for a straightforward grid search over a limited amount of precision pairs such as K8V4, K4V8, K4V2, and so on. Therefore, this allocation split of bits is not the core focus of the framework, with the focus instead being on deploying a layerwise quantization scheme.
Two other studies also touch upon the key-value difference, but like KVTuner, do not provide much depth into the difference. One of these studies is a KV quantization framework called SKVQ, which deploys a sliding window filter along the input tokens to remove tokens that fall outside the window frame and that meet certain criteria [33]. The quantization strategy achieves 2 bit key quantization and 1.5 bit value quantization, but the authors do not explore the possible reasons behind why this difference in bit allocation yielded such powerful results, instead arriving at this bit allocation via hyperparameter search.
The other study is a KV quantization framework called QAQ, which uses the attention output to perform the key quantization and the self-attention output to perform the value quantization [32]. Unlike SKVQ and all other KV quantization strategies, QAQ is the first to target the keys and values in different approaches. However, the difference in the approach lies in the type of data being used to perform the quantization, not in the exact number of bits being allocated to one versus the other. Finally, both SKVQ and QAQ are post-training quantization approaches, making them susceptible to overhead issues during the attention computation.
Figure 2: Frobenius norm for Llama Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The norms are higher for the Key weights than for the Value weights across layers.
3 Dynamics of the Norms of Key and Value
3.1 Key-Value Norm Disparity
Before formalizing our finding, a brief examination is presented on how the magnitudes of the key and value projection matrices influence KV cache quantization. In the self-attention mechanism, the key weights $W^{k}$ scale the token representations that enter the cache, whereas the value weights $W^{v}$ scale the representations that are read out once the attention scores are formed. Because quantization noise is approximately proportional to the dynamic range of the signal, the relative sizes of $\|W^{k}\|$ and $\|W^{v}\|$ determine which part of the cache is more sensitive to a given bit budget. Computing norms across four Llama-3 models (Fig. 2) show that the Frobenius norm of $W^{k}$ consistently exceeds that of $W^{v}$ in almost every layer; the same ordering holds for the spectral norm. Intuitively, this systematic gap motivates the formal statement captured in Theorem Key-Value Norm Disparity below.
Theorem 1 (Key-Value Norm Disparity).
The expectation of the Frobenius norm for the key weight matrices are greater than that for the value weight matrices.
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (1)
The proof of Theorem 1 is deferred to Appendix B, but its core argument can be sketched as follows: Starting from a standard Xavier initialization[36], the analysis tracks how stochastic-gradient descent updates the Frobenius norms of $W^{k}$ and $W^{v}$. Under isotropic-input assumptions, the cross-term $\langle W^{m},\,\partial\mathcal{L}/\partial W^{m}\rangle_{F}$ becomes negligible, so the expected norm growth is governed by the squared gradient magnitudes. A careful decomposition of those gradients shows that the Key weight update contains an additional factor of the query weight matrix $Q$, whose own norm increases during training. Consequently, the expected squared gradient of $W^{k}$ dominates that of $W^{v}$, causing $\mathbb{E}\bigl{[}\|W^{k}\|_{F}\bigr{]}$ (and, by monotonicity, the spectral norm) to outpace its value counterpart. This asymptotic divergence establishes the persistent norm gap asserted Key-Value Norm Disparity.
Although Fig. 2 illustrates this disparity for the Llama family, supplementary plots for Mistral models in Appendix D confirm that the same pattern holds across diverse architectures [8]. The next subsection investigates how this norm disparity influences quantization error, laying the groundwork for data-driven bit-allocation strategies.
3.2 Key-Driven Quantization
Theorem Key-Value Norm Disparity establishes a persistent norm gap between the Key and Value weight matrices. Since quantization noise scales with the dynamic range of the signal, weight matrices of larger norm inevitably introduce larger absolute errors under identical bit precision. Consequently, assigning equal quantization budgets to the Key and Value caches is sub-optimal.

### v3

# Quantize What Counts: More for Keys, Less for Values

@@ Abstract
Large Language Models (LLMs) suffer inference-time memory bottlenecks dominated by the attention Key-Value (KV) cache, which scales with model size and context length. While KV-cache quantization alleviates this cost, bit allocation between keys and values is often tuned heuristically, lacking theoretical grounding and generalizability. This paper proposes two theorems that anchor mixed-precision KV quantization in the intrinsic geometry of Transformer models. First, key projections systematically have larger spectral and Frobenius norms than value matrices, implying higher information density along the key path. Second, for any given memory budget, prioritizing precision for keys over values strictly reduces quantization error and better preserves accuracy. Empirical evaluations across various prominent LLMs and benchmarks show that key-favored allocations (e.g., 4-bit keys, 2-bit values) retain up to 98.3% accuracy compared to uniform allocations (e.g., 4-bit for both), while conserving memory. These results transform bit allocation from ad hoc tuning into a theoretically grounded, geometry-driven design principle for efficient LLM inference. Source code is available at https://github.com/mohsenhariri/spectral-kv.
Quantize What Counts: More for Keys, Less for Values
Mohsen Hariri1, Alan Luo1, Weicong Chen1, Shaochen Zhong2, Tianyi Zhang2, Qifan Wang3, Xia Hu2, Xiaotian Han1, Vipin Chaudhary1 1Case Western Reserve University; 2Rice University; 3Meta mohsen.hariri@case.edu
@@ 1 Introduction
Large Language Models (LLMs) have rapidly scaled in recent years, driving major advances in generative capabilities and reasoning performance Vaswani et al. (2017); Sutskever et al. (2014). Model size has increased by several orders of magnitude: GPT has grown from 117M parameters in GPT-1 Radford et al. (2018) to 1.5B in GPT-2 Radford et al. (2019), 175B in GPT-3 Brown et al. (2020), and 1.8T in GPT-4 OpenAI et al. (2024). Open-source models have followed a similar trajectory, with Llama reaching 2T parameters Meta AI (2025b), Mistral Large scaling to 123B Mistral AI (2024), and DeepSeek V3 to 671B DeepSeek-AI et al. (2025b).
@@
70B 0.76 0.87
Figure 1: Key cache needs more bits. (Left): Spectral norms of the key cache (blue) and value cache (orange) across layers in Llama3.3-70B show that key caches consistently exhibit higher norms. (Right): GSM8k accuracy for two schemes: $\text{K}_{2}\text{V}_{4}$, representing 2-bit allocation for the K cache and 4-bit allocation for the V cache and $\text{K}_{4}\text{V}_{2}$, representing 4-bit allocation for the K cache and 2-bit allocation for the V cache, demonstrates that allocating more bits to the key cache maintains strong performance, confirming the efficacy of norm-aware, mixed-precision quantization.
However, this rapid growth has introduced severe inference-time memory bottlenecks, primarily due to the Key-Value (KV) cache Wei et al. (2022). As parameter counts increase, context lengths must also grow to support more complex reasoning, which further expands the KV cache and strains GPU memory Sheng et al. (2023a). Modern systems already support extremely long contexts Google DeepMind (2025); OpenAI (2024, 2025); DeepSeek-AI et al. (2025a), reaching up to 10 million tokens Meta AI (2025a), making memory-efficient methods essential.
KV cache quantization refers to reducing the precision of KV tensors (e.g., BF16 to INT4), which can provide substantial memory savings while maintaining controlled accuracy degradation, provided applied strategically Han et al. (2016); Nagel et al. (2021). Although many KV quantization methods have been proposed, most determine key–value bit splits through ad hoc hyperparameter tuning on cache statistics (e.g., inference-time activations
) rather than grounding them in intrinsic model properties (e.g., model weights). This raises a fundamental question: How should bits be allocated in a principled and generalizable way?
To answer, Figure 1 illustrates two key observations. First, key caches ($\mathbf{K}$) consistently have larger spectral norms than value caches ($\mathbf{V}$). Second, assigning more bits to keys (e.g., $\mathrm{K}_{4}\mathrm{V}_{2}$) improves accuracy across multiple models compared to key-underprovisioned allocations ($\mathrm{K}_{2}\mathrm{V}_{4}$). These motivate a deeper investigation into the distinct roles of key-value weights (i.e., $\mathbf{W}^{\mathbf{K}}$ and $\mathbf{W}^{\mathbf{V}}$) in attention mechanisms and their implications for quantization performance. Toward this, our contributions are:
  • We propose the Key-Value Norm Disparity theorem, proving that the expected spectral and Frobenius norms of $\mathrm{W}^{\mathrm{K}}$ predominantly exceed those of $\mathrm{W}^{\mathrm{V}}$ across prominent LLM models (i.e., Llama3 and Mistral herds). We then derive the Key-Prioritized Quantization theorem, establishing the theoretical foundation on why assigning higher precision to $\mathrm{K}$ than $\mathrm{V}$ strictly reduces quantization error, enabling greater KV-cache compression while maintaining accuracy.
  • We corroborate and operationalize these theorems across a diverse set of models (Llama-3.2-1B/3B/8B, Llama-3.3-70B, Phi-4-14B, Qwen3-0.6B/1.7B/4B/8B, DeepSeek-R1, Mistral-0.3-7B), datasets (C4, MMLU, GSM8K, EQ-Bench, CoQA, and LongBench [^1] We purposefully select generative tasks rather than commonsense reasoning tasks to better isolate quantization effects; see Section C.1 for details.), and two quantization backends (Optimum Quanto and HQQ).
    Notably, $\mathrm{K}_{4}\mathrm{V}_{2}$ retains 98.3% (1-shot) and 94.1% (5-shot) accuracy of $\mathrm{K}_{4}\mathrm{V}_{4}$ accuracy while reducing KV-cache memory by 25%, demonstrating both the theoretical soundness and practical effectiveness of the proposed strategy.
  • Owing to its efficient one-off tunability, we show that our geometry-driven mixed-precision strategy is orthogonal to existing inference-time KV quantization methods and can be seamlessly integrated to yield synergistic gains. In a case study with rotation-based outlier redistribution, combining a key-prioritized quantization ($\mathrm{K}_{4}\mathrm{V}_{2}$) with key-only rotation outperforms $\mathrm{K}_{4}\mathrm{V}_{4}$ by 4.4-18% in accuracy across tasks. In contrast, rotating value caches is unnecessary and sometimes detrimental
2 Background and Related Work
KV Quantization.
Quantization methods for LLMs can be categorized by timing: training-time and post-training (PTQ) Gholami et al. (2021). Training-time quantization integrates quantization into model training, typically achieving higher accuracy by quantizing weights or activations during the optimization process. However, it requires labeled data and incurs significant training overhead. PTQ applies quantization after training, avoiding retraining costs and labeled data requirements Nagel et al. (2021), but sometimes yields lower accuracy.
PTQ can target different model components: weights, activations, or the KV cache. Weight-only quantization Frantar et al. (2023a, b); Lin et al. (2024) achieves strong accuracy but does not reduce activation or KV memory. Weight-activation quantization Dettmers et al. (2022); Xiao et al. (2023); Shao et al. (2024) reduces overall memory but often sacrifices accuracy. KV quantization offers the best of both: it targets the rapidly growing KV cache Pope et al. (2023), providing activation-level memory savings while maintaining the accuracy of weight-only methods Yue et al. (2024).
Existing KV Quantization Schemes.
KV quantization methods can be categorized by how they treat keys and values Li et al. (2025a):
  • Outlier redistribution. These methods smooth or relocate outliers (i.e., unusually large activation values that dominate quantization ranges) in KV tensors, e.g., SmoothQuant Xiao et al. (2023), AWQ Lin et al. (2024), and OmniQuant Shao et al. (2024).
  • Fixed-precision. A single bit-width is used for both keys and values, ignoring their different roles and statistical properties Yao et al. (2022); Sheng et al. (2023b); Liu et al. (2024).
  • Mixed-precision. Different bit-widths are assigned to different parts of the cache Hooper et al. (2024); Yue et al. (2024); Li et al. (2024); Dong et al. (2024); Duanmu et al. (2024); Zhang et al. (2024); Li et al. (2025b).
While mixed-precision schemes are the most flexible, existing methods have not systematically explored asymmetric bit allocation between keys and values. KVQuant Hooper et al. (2024), for example, focuses on vector-wise outlier handling rather than analyzing keys and values separately. Only a few methods have attempted to address this issue, and even then, only superficial insights have been gained.
Needs and Gaps.
Despite rapid progress in KV quantization, several needs from the LLM research community remain unmet. First, there is a need for principled strategies to guide bit allocation. Current frameworks either treat the key-value split as a hyperparameter tuned through grid search Li et al. (2025b) or rely on heuristics derived from cache statistics (e.g., activation ranges or distributions collected at inference time) Zhang et al. (2023); Liu et al. (2025); Yang et al. (2024). These approaches are costly, model- or data-specific, and provide little theoretical insight into the inherent differences between keys and values. Second, there is a need to understand and exploit key-value asymmetry. Existing works such as KVTuner Li et al. (2025b), SKVQ Duanmu et al. (2024), and QAQ Dong et al. (2024) briefly observe that allocating more bits to keys can preserve accuracy, but none explain why or propose a generalizable strategy. KVTuner reports differences in attention vs. perplexity errors across bit pairs without analysis; SKVQ finds asymmetric allocations (e.g., 2-bit keys, 1.5-bit values) through hyperparameter search rather than model structure; and QAQ focuses on different data types for keys and values rather than bit-width asymmetry. Third, there is a need for lightweight, modular methods that integrate seamlessly with existing KV quantization frameworks. The studies mentioned above often involve complex, runtime-dependent procedures that are difficult to generalize or combine, limiting their drop-in applicability and composability with other techniques.
Our Perspective.
We address these needs by deriving bit allocation strategies directly from model weights, analyzing spectral and Frobenius norms of key and value matrices. Our approach is lightweight, incurring only a one-off analytical cost per model without any inference-time introspection; generalizable, since weight statistics are invariant across inputs and tasks; and principled, grounding allocation in linear algebraic properties rather than heuristic search. Moreover, our mix-precision quantization oracle is orthogonal to existing KV quantization techniques, paving a foundation that others can build upon. For example, pairing our mixed-precision allocation with rotation-based outlier redistribution techniques yields complementary improvements in both accuracy and memory savings. In this way, we elevate bit allocation from an ad hoc design choice to a geometry-informed building block for future KV quantization frameworks.
3 Norm Dynamics of KV Weights
We now establish a theoretical foundation for mixed-precision KV-cache quantization by analyzing the intrinsic geometry of the key and value projection weights. Our analysis proceeds in two steps. First, we prove that key weights systematically exhibit larger spectral and Frobenius norms than value weights (Key-Value Norm Disparity), a property that is preserved in the resulting key and value caches. Second, we demonstrate that this norm gap directly impacts quantization error, providing a formal guarantee that assigning higher bit precision to keys yields strictly lower distortion and higher inference accuracy than symmetric or inverted allocations (Key-Prioritized Quantization).
3.1 Key-Value Norm Disparity: Key Weights Dominate in Norm
The key weight matrix $\mathrm{W}^{K}$ maps hidden states into the key cache, whereas the value weight matrix $\mathrm{W}^{V}$ determines the representations retrieved during attention. Because quantization error scales approximately with the dynamic range of the signal, the relative magnitudes of $\|\mathrm{W}^{K}\|$ and $\|\mathrm{W}^{V}\|$, e.g., their Frobenius or spectral norms, indicate which cache is more sensitive to quantization.
Figure 2: Frobenius norms of key and value weight matrices across the Llama 3 family. $\|W^{K}\|_{F}$ consistently exceeds $\|W^{V}\|_{F}$ across nearly all layers, with the exception occurring in early layers of the 70B variant.
Empirical measurements across four Llama-3 herds (Figure 2) show that the Frobenius norm of $\mathrm{W}^{K}$ consistently exceeds that of $\mathrm{W}^{V}$ in nearly every layer; the same ordering holds for spectral norms. This persistent gap motivates the following theorem.
Theorem 1 (Key-Value Norm Disparity) Let $\mathrm{W}^{K}$ and $\mathrm{W}^{V}$ denote the key and value projection matrices in a Transformer. Then $\mathbb{E}\bigl[\|\mathrm{W}^{K}\|_{F}\bigr]>\mathbb{E}\bigl[\|\mathrm{W}^{V}\|_{F}\bigr].$
The detailed proof is provided in Appendix A.2. The core idea here is to examine how the Frobenius norms of the key and value weight matrices evolve during training. The analysis begins with Xavier initialization Glorot and Bengio (2010), where all projection matrices have identical expected norms, and tracks their evolution under stochastic gradient descent (SGD). Intuitively, $\mathrm{W}^{K}$ play a dual role: they shape the attention map and determine what representations are stored in the cache. Each input is multiplied by $\mathrm{W}^{K}$ to produce keys that interact multiplicatively with the queries derived from $\mathrm{W}^{Q}$ (i.e., the query projection). As training proceeds, $\mathrm{W}^{Q}$ typically grows to sharpen attention, amplifying the gradient signals backpropagated into $\mathrm{W}^{K}$. By contrast, $\mathrm{W}^{V}$ only influences post-attention representations, so its gradients lack this multiplicative amplification. This architectural asymmetry causes $\mathrm{W}^{K}$ to receive systematically larger updates, leading to persistently larger norms over time.
This phenomenon is ubiquitous. Appendix C.2 offers analogous results for Mistral family Mistral AI (2024), demonstrating the generality of the pattern. We next show that this norm disparity has direct consequences for quantization.
3.2 Key-Prioritized Quantization: Key-Favored Allocation Minimizes Quantization Error
Theorem 3.1 implies that, on average, $\mathrm{W}^{K}$ and their resulting activations, $\mathrm{K}$, have larger magnitude than their value counterparts. Since quantization error under uniform scalar quantization scales with the signal energy, assigning equal bit precision to both is sub-optimal.